A solid cylinder of radius r1 = 2.2 cm, length h1 = 5.9 cm, emissivity 0.91, and temperature 38°C is suspended in an environment of temperature 70°C. (a) What is the cylinder's net thermal radiation transfer rate P1? (b) If the cylinder is stretched until its radius is r2 = 0.55 cm, its net thermal radiation transfer rate becomes P2. What is the ratio P2/P1?
Answer #1

a)

r1 = radius = 2.2 cm = 0.022 m

h1 = height = 5.9 cm = 0.059 m

A1 = Area of cylinder = 2\pir1 (r1 + h1) = (2 x 3.14 x 0.022) (0.022 + 0.059) = 0.011 m2

T = temp. of surrounding = 70 C = 70 + 273 K = 343 K

To = cylinder's temp. = 38 C = 38 + 273 = 311 K

P1 = \epsilon\sigma A1 (T4 - To4)

P1 = (0.91) (5.67 x 10-8) (0.011) (3434 - 3114)

P1 = 2.55 Watt

b)

initial Volume of cylinder = final volume of cylinder after being stretched

\pir12 h1 = \pir22 h2

(0.022)2 (0.059) = (0.0055)2 h2

h2 = 0.944 m

A2 = Area of cylinder = 2\pir2 (r2 + h2) = (2 x 3.14 x 0.0055) (0.0055 + 0.944) = 0.033 m2

P2 = \epsilon\sigma A2 (T4 - To4)

P2 = (0.91) (5.67 x 10-8) (0.033) (3434 - 3114)

P2 = 7.64 Watt

Ratio : P2/P1 = 7.64/2.55 = 3